# GeoTransformer: Enhancing Urban Forecasting with Geospatial Attention Mechanisms

### v1

@@ Abstract
Recent advancements have focused on encoding urban spatial information into high-dimensional spaces, with notable efforts dedicated to integrating sociodemographic data and satellite imagery. These efforts have established foundational models in this field. However, the effective utilization of these spatial representations for urban forecasting applications remains under-explored. To address this gap, we introduce GeoTransformer, a novel structure that synergizes the Transformer architecture with geospatial statistics prior. GeoTransformer employs an innovative geospatial attention mechanism to incorporate extensive urban information and spatial dependencies into a unified predictive model. Specifically, we compute geospatial weighted attention scores between the target region and surrounding regions and leverage the integrated urban information for predictions. Extensive experiments on GDP and ride-share demand prediction tasks demonstrate that GeoTransformer significantly outperforms existing baseline models, showcasing its potential to enhance urban forecasting tasks.
@@ Introduction
In urban forecasting tasks involving predicting economic indicators and human mobility, classical methods usually rely on low-dimensional numeric data, which includes Point of Interest (POI) data(Li et al. 2022), survey data(Roskam and Kunst 2008; DHS et al. 2013), GPS records (Wang et al. 2023a; Moreira-Matias et al. 2013; Pappalardo et al. 2013), demographic census, geospatial features and so forth.
Recently, many researchers have explored using high-dimensional information to fully capture the complexities of urban dynamics. A main direction is utilizing urban imagery for feature extraction and prediction (Ayush et al. 2020; Li et al. 2022; Yeh et al. 2020; Abdelhalim and Zhao 2024). A more advanced direction is encoding the urban information to a high-dimensional space representation, either through spatial representation learning (Mai, Li, and Lao 2023; Wu et al. 2024) or fusing satellite images and sociodemographic data (Wang et al. 2024). While these studies have established frameworks for encoding urban information, there has been limited investigation into optimizing these high-dimensional data for urban forecasting.
Numerous established methods in natural language processing (NLP) and computer vision have effectively harnessed high-dimensional data. Particularly, Transformer architecture and attention mechanisms have proven highly effective in decoding high-dimensional representations of natural language (Vaswani et al. 2017a; Devlin et al. 2018; Radford et al. 2018) and images (Dosovitskiy et al. 2020a; Rombach et al. 2022). Urban representations, characterized by high dimensionality and spatial correlations, are well-suited for applying similar decoding methods. Some studies have explored the use of attention mechanisms to capture spatial dependencies, particularly through Graph Attention Networks (GAT) (Feng and Tassiulas 2022; Zheng et al. 2020). However, these approaches primarily depend on graph structures and spatial embeddings. In fact, incorporating geospatial statistical priors might be more effective in encoding spatial dependencies, potentially reducing the need for complex data structuring and simplifying the system design.
To address this issue, we propose GeoTransformer, a deep architecture akin to Transformer (Vaswani et al. 2017b) that combines geospatial characteristics prior to the attention mechanism. The novel geospatial attention mechanism calculates attention between different urban regions weighted by distance, incorporating information from surrounding areas. This approach implicitly includes spatial dependencies in the attention mechanism. Extensive experiments demonstrate that our model significantly outperforms several baseline methods in predicting GDP and ride-share demand.
@@
  • We propose GeoTransformer, a transformer-akin deep architecture designed for decoding high-dimensional urban representations and enhancing prediction accuracy.
  • We introduce a geospatial attention mechanism that prioritizes attention between regions weighted by spatial proximity, assigning higher weights to closer regions.
@@
High Dimensional Input for Urban Forecasting
Urban forecasting leverages multiple data sources and advanced computational techniques to address challenges faced by cities. Traditional methods usually adopt low-dimensional numeric data for prediction. In recent years, with the need rising for complex prediction tasks and the development of deep learning techniques, researchers have explored various ways to leverage high-dimensional urban information. One direction is directly using various urban imagery for feature extraction and prediction. Albert, Kaur, and Gonzalez employ Convolutional Neural Networks (CNNs) to detect land-use patterns from satellite images. Ayush et al.; Li et al.; Yeh et al. utilize urban imagery to predict socioeconomic status. Abdelhalim and Zhao leverage roadside imagery to predict transit travel time.
Recently, encoding urban information into a high-dimensional representation marks a new direction in enhancing flexibility and prediction accuracy for different urban forecasting tasks. Mai, Li, and Lao develop spatial representation learning (SRL), utilizing deep neural networks (DNNs) to encode and feature spatial data in the forms of points, polylines, polygons, graphs, etc. Wang et al. propose a theoretical framework deep hybrid model (DHM) fusing demographic numeric data and satellite images into urban latent representations. However, how to leverage these urban representations more effectively in prediction tasks remains unexplored.
Transformer and attention mechanism(Vaswani 2017) has been widely adopted for its efficiency in capturing relationships within high-dimensional representations. They are first proposed to solve natural language understanding and then visual models like ViT(Dosovitskiy et al. 2020b) and Stable Diffusion (Rombach et al. 2022) also utilize attention mechanisms to comprehend images and the spatial relationships between image patches. Given its strength in handling high-dimensional data and capturing spatial relationships, the attention mechanism is well-suited for adaptation in modeling urban high-dimensional representations.
Approaches to Modeling Spatial Dependency
@@
Spatial dependency has traditionally been addressed using two primary approaches. First, feature engineering methods like k-nearest neighbors (KNN) are employed to convert surrounding spatial information into features. Second, spatial statistical methods such as spatial error regression, spatial lag models(Anselin 2013), and geographically weighted regression (GWR) (Fotheringham, Charlton, and Brunsdon 1998) have been widely used to account for spatial autocorrelation.
In recent years, Graph Neural Networks (GNNs) have become a widely adopted method for capturing spatial dependencies, especially in spatiotemporal prediction(Wang et al. 2023b; Zhuang et al. 2023; Cai et al. 2020). Urban built environment data such as road networks, are inherently suitable for representation as graph structures due to their topological properties. However, grid-based high-dimensional urban representations often lack clear graph-like structures, making graph construction particularly challenging. Moreover, the intrinsic complexity of such high-dimensional data can further hinder the effective training of models.
Recent advancements have seen the integration of graph structures with attention mechanisms (Zheng et al. 2020; Feng and Tassiulas 2022), particularly through Graph Attention Networks (GATs)(Zhang, Yu, and Liu 2019). While these models incorporate spatiotemporal attention mechanisms, their reliance on graph structures and spatial embeddings remains significant for capturing spatial dependencies. Integrating spatial statistical priors could simplify the model’s reliance on complex data structures and streamline system design. Tobler’s First Law of Geography (Tobler 1970), which posits that closer entities are more likely to be related, could provide a valuable framework for capturing spatial dependency. Despite its potential, the integration of spatial statistical priors with attention mechanisms remains unexplored.
Therefore, in this study, we aim to introduce an attention mechanism guided by spatial statistical priors as a new paradigm for effectively leveraging high-dimensional urban representation data.
@@
Gross Domestic Product (GDP) is a critical indicator in urban studies, offering vital insights into the economic health, developmental patterns, and sustainability of cities(Ribeiro et al. 2021; Mahtta et al. 2022; Brilhante and Klaas 2018). Leveraging Nighttime Light (NTL) data to estimate GDP as an economic indicator has become a widely adopted approach (Liu et al. 2021; Sun et al. 2020). There is substantial research indicating a strong association between GDP and urban spatial characteristics (Zhang, He, and Yuan 2020; Zhaoa and Chenb 2023), with numerous studies employing machine learning methods to model and predict regional GDP using structured urban information or urban imagery data(Kopczewska 2022; Chen et al. 2022b). For example, Chen et al. utilize deep learning methods to predict county-level GDP grids. While substantial research has established the relationship between GDP and urban spatial characteristics, the application of high-dimensional urban representations for GDP prediction remains underexplored. Given the critical role of GDP in validating urban prediction models, we selected GDP estimates derived from Nighttime Light (NTL) data as a benchmark for evaluating our approach.
In addition to economic indicators, mobility indicators such as ride-share demand serve as a critical indicator for urban forecasting, and it has significant implications for transportation planning, economic activity, and population behavior (Komanduri et al. 2018; Ma and Hanrahan 2020; Zhang and Zhang 2018). For instance, Shaheen et al. emphasize that ride-share services can influence traffic patterns, reduce private vehicle ownership, and improve accessibility in urban environments. Researchers have developed utilizing deep learning methods to predict ride-share demand from urban spatial information(Mühlematter et al. 2023; Zhao et al. 2022; Zhang and Zhao 2022). Zhao et al. couple graph neural networks (GNNs) with the spatial-temporal influence of the built environment. Zhang and Zhao propose a Clustering-aided Ensemble Model to improve ridesourcing demand forecasting. Similarly, few studies have explored using high-dimensional urban representations to predict ride-share demand. Given the significance of ride-share demand as a key indicator with strong spatial dependencies, it serves as an ideal benchmark for evaluating the efficacy of our proposed method.
@@
As illustrated in Figure 1, our framework adopts an encoder-decoder structure. The encoder leverages existing state-of-the-art methods to process numeric and imagery inputs, representing city regions in a latent space through a mixing operator. Our novel GeoTransformer decoder model then utilizes this latent representation to predict GDP and ride-share demand.
@@
Encode the city: Mixing Operator
To fuse satellite imagery and sociodemographic information into a high-dimensional latent representation, we adopt a mixing operator $\mathcal{M}$, which can be represented as:
$z_{n}=\mathcal{M}(x_{n},I_{n})$ (1)
in which $x_{n}$ and $I_{n}$ represent the sociodemographics and satellite imagery inputs, and $z_{n}$ represents the latent representation of an urban area.
Following the work Deep Hybrid Model (Wang et al. 2024), we adopt a supervised autoencoder (SAE). The SAE maps the imagery into latent variables through the encoder $E(I_{n})$, and reconstructs images through the decoder $D(z_{n})$ and sociodemographic through a supervision network $F(z_{n})$ (Figure 2 ). It is trained by optimizing the sum of autoencoder loss $\mathcal{L}_{AE}$ and sociodemographic supervision loss $\mathcal{L}_{sup}$:
$\mathcal{L}_{SAE}=(1-\lambda)\mathcal{L}_{AE}+\lambda\mathcal{L}_{sup}$ (2)
where $\mathcal{L}_{sup}$ denotes the absolute error of the sociodemographic variables $x$ $\mathcal{L}_{AE}$ represents the sum of L1 image reconstruction loss $\mathcal{L}_{rec}$, the KL divergence loss $\mathcal{L}_{kl}$, and the learned perceptual image patch similarity loss $\mathcal{L}_{lpipis}$:
$\mathcal{L}_{sup}=\alpha_{0}|x-\hat{x}|$ (3)
$\mathcal{L}_{AE}=\alpha_{1}\mathcal{L}_{rec}+\alpha_{2}\mathcal{L}_{kl}+\alpha_{3}\mathcal{L}_{lpipis}$ (4)
in which all the $\alpha$ terms are the adaptive hyperparameters used for scaling each loss.
For each autoencoder loss, $\mathcal{L}_{rec}$ measures the L1 distance between the input images $I_{n}$ and reconstructed images $\hat{I_{n}}$ in the pixel space, $\mathcal{L}_{kl}$ measures the distance between the estimated latent distribution and the standard normal prior distribution, and $\mathcal{L}_{lpips}$ is calculated by extracting feature stacks from $L$ layers of a pre-trained neural network, unit-normalizing them in the channel dimension ($\hat{y}^{l}_{I_{n,hw}},\hat{y}^{l}_{\hat{I}_{n,hw}}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$), and scaling by vector $w_{l}$ before computing the L2 distance.
The mixing operator allows for the fusion of sociodemographics and satellite imagery information in a high-dimensional latent space. Each region of a city is then encoded into a latent representation using the encoder of the trained SAE.
@@
GeoTransformer: Geospatial-weighted Attention Mechanisms
We propose a novel approach GeoTransformer to decode the urban latent representations for downstream tasks including GDP and ride-share demand prediction.
As shown in Figure 3, our method adopts a structure akin to the Transformer (Vaswani 2017) decoder. Since the downstream tasks do not involve sequence prediction, the self-attention module is optional in this study. We introduce a geospatial attention mechanism to capture the spatial dependencies and provide location information between city regions. The outputs of the decoder layers are then passed through a fully connected linear layer. The whole system $\mathcal{G}$ can be represented as:
$y_{n}=\mathcal{G}(z_{n},\text{kNN}(z_{n},Z))$ (5)
in which the input includes the latent representation of a region $z_{n}$ and its k-nearest neighbors’ latent representations $\text{kNN}(z_{n},Z)$ from all regions $Z$, and $y_{n}$ is the predicted output.
The geospatial attention mechanism is the core component of our method. It computes attention scores between city regions and applies geospatial weighting to these scores. The attention calculation is represented as:
$\text{GeoAtt}(Q,K,V)=W\odot\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (6)
@@
where the query matrix $Q$ contains the latent representation of the query region $z_{n}$ for which the target variable is being predicted, while the value matrix $V$ consists of $z_{n}$ itself and the latent representations of its neighbors as value regions. $d_{k}$ denotes the dimension of the key vectors. Intuitively, as the similarity between two city regions is not directly related to target variables, we set the key matrix $K$ as trainable weights rather than being identical to $V$.
@@
$W$ is the spatial weighting factor, computed based on the spatial distance between query regions in $Q$ and value regions in $V$. Our model provides linear and non-linear spatial weighting options based on the assumption of Tobler’s First Law of Geography (Tobler 1970), suggesting that regions closer to the query region are assigned higher weights.
Linear Normalization
scales the distances within 0 to 1, inversely transforming shorter distances into higher weights:
$W_{j}=1-\frac{d_{j}}{\text{max}(d)}$ (7)
@@
Inverse Distance Weighting (IDW) (Shepard 1968)
calculates weights inversely proportional to the distance:
$W_{j}=\frac{1}{d^{p}_{j}+\epsilon}$ (8)
where $d_{j}$ is the distance to the $j^{th}$ value region, the power value $p$ is set to 2 in standard applications, and $\epsilon$ is a small constant to prevent division by zero. The weights are then subsequently normalized to ensure they sum to unity.
Gaussian Weighting
applies a Gaussian function to the distances, providing a gradual decrease in weight as distance increases:
$W_{j}=e^{-(\frac{d^{2}_{j}}{2\sigma^{2}})}$ (9)
where $d_{j}$ is the distance to the $j^{th}$ value region, and $\sigma$ determines the spread of the weights. The weights are then subsequently normalized to ensure they sum to unity.
Adhering to the established multi-head attention mechanism (Vaswani 2017), our model employs multiple attention heads to capture various spatial relationships concurrently. The calculation can be represented as:
@@
where $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$ and $W^{O}$ are the projection matrices. The amount of heads and layers of the decoder module is also adaptive.
The GeoTransformer architecture is designed to extract geospatial information in high-dimensional latent space, and it is also adaptable to various geospatial prediction tasks due to the flexibility of the output layers. The design of geospatial attention allows the model to capture and integrate spatial dependencies, utilizing local context from both the target region and its surroundings for more accurate predictions. Leveraging high-dimensional urban information and deep architectures, our GeoTransformer represents a novel approach to geospatial analysis and urban planning applications.
Figure 3: The GeoTransformer structure includes a geospatial attention module that calculates the geospatial weighted attention scores between the query region and its neighboring value regions.
Panel 1 Baselines with only numeric or imagery input
Model 1: Linear Regression sociodemographic numeric input (12 dim)
Model 2: ViT satellite imagery input (4*512*512 dim), heads = 12, layers = 12
Panel 2 Baselines with urban latent representations input
Model 3: Polynomial Regression latent input (reduced dim from 4096 to 128 using PCA), poly degree = 2, $l_{2}$ regularization with $\lambda$ = 0.5
Model 4: GAT latent input (4096 dim), layers = 4
Panel 3 GeoTransformer with 3 geospatial weighting methods
Model 5: Linear Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
Model 6: IDW Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
Model 7: Gaussian Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=81 / 169
@@
Panel 1 Baselines with only numeric or imagery input
1 Linear Regression 3.22/3.13 1.44/1.43 0.159/0.151 8.45/6.11 8.36/5.99 0.668/0.720
2 ViT 0.616/0.828 0.595/0.703 0.822/0.759 15.6/22.5 4.10/4.16 0.526/0.315
Panel 2 Baselines with urban latent representations input
3 Polynomial Regression 2.19/3.25 1.17/1.46 0.429/0.118 12.8/19.5 5.05/6.00 0.494/0.335
4 GAT 0.469/0.726 0.476/0.602 0.866/0.805 7.21/9.51 2.01/2.35 0.750/0.575
Panel 3 GeoTransformer with 3 geospatial weighting methods
5 Linear Weighting 0.081/0.216 0.210/0.346 0.979/0.944 1.62/7.26 0.98/2.53 0.934/0.779
6 IDW Weighting 0.570/0.686 0.537/0.585 0.849/0.822 0.172/3.02 0.491/1.16 0.993/0.908
7 Gaussian Weighting 0.448/0.567 0.486/0.533 0.882/0.852 0.263/9.78 0.469/3.97 0.989/0.703
@@
In this section, we provide our method’s results in GDP and ride-share demand prediction and compare them with baseline methods. Ablation experiments are conducted to demonstrate the effectiveness of the geospatial attention component and model architecture design.

### v2

# GeoTransformer: Enhancing Urban Forecasting with Dependency Retrieval and Geospatial Attention

@@ Abstract
Recent advances in urban forecasting have leveraged high-dimensional spatial data through two primary approaches: graph-based methods that rely on predefined spatial structures and region-based methods that use satellite imagery for local features. Although these methods have laid an important foundation, they struggle to integrate holistic urban information and dynamically model spatial dependencies. To address this gap, we propose GeoTransformer, a framework combining high-dimensional regional embeddings with dynamic spatial modeling. GeoTransformer features two innovations: (1) a dependency retrieval module identifying spatial dependencies to select relevant regions, and (2) a geospatial attention mechanism leveraging global urban information. These components unify structural and global urban information for better predictions. Extensive experiments on GDP and ride-share demand forecasting show that GeoTransformer outperforms baselines, highlighting its effectiveness in advancing urban forecasting tasks.
@@ Introduction
In urban forecasting tasks involving predicting economic indicators and human mobility, classical methods usually rely on low-dimensional numeric data, which includes spatial features, socioeconomic data, human mobility data, and so forth (Li et al. 2022; Roskam and Kunst 2008; DHS et al. 2013; Wang et al. 2023; Moreira-Matias et al. 2013; Pappalardo et al. 2013).
Recently, advances in natural language processing (NLP) and deep learning have enabled the use of high-dimensional information to capture the complexities of urban dynamics. Some studies have transformed urban information into high-dimensional representations, either through text embedding (Chen et al. 2022b; Li et al. 2020, 2023), spatial representation learning (Mai, Li, and Lao 2023; Wu et al. 2024b) or by encoding satellite imagery data (Jean et al. 2018; Wang et al. 2024), offering a new urban modeling paradigm by leveraging high-dimensional data to capture complex urban dynamics.
High-dimensional urban forecasting applications can be broadly categorized into two directions. The first utilizes graph-based modeling with spatial feature embeddings, then using Graph Neural Networks (GNNs) or Graph Attention Networks (GATs) for predictions (Feng and Tassiulas 2022; Zheng et al. 2020; Chen et al. 2022b; Li et al. 2020, 2023). While effective, these methods depend heavily on predefined spatial structures and high-quality spatial data, which limits their flexibility in data-sparse or dynamically changing environments.
Another direction focuses on deriving high-dimensional representations directly from satellite imagery (Jean et al. 2018; Wang et al. 2024). These methods avoid reliance on predefined structures by encoding urban regional information into independent high-dimensional patches. However, these approaches only utilize local information within each patch for prediction and lack the capability to incorporate global urban context (Wang et al. 2024), which is crucial for tasks requiring holistic understanding.
To address these limitations, we propose GeoTransformer, a novel transformer-based (Vaswani et al. 2017) architecture that leverages high-dimensional regional embeddings without requiring predefined spatial structures. GeoTransformer incorporates two key innovations: (1) a dependency retrieval module that infers and integrates urban structural dependencies to retrieve the most relevant regions; and (2) a geospatial attention mechanism, utilizing cross-attention to dynamically capture global urban information, enhanced with spatial proximity and information entropy weighting. These innovations enable GeoTransformer to effectively combine local and global perspectives, offering a robust framework for diverse urban prediction tasks.
Extensive experiments demonstrate that our model significantly outperforms baseline methods in predicting GDP and ride-share demand.
@@
  • We propose GeoTransformer, a novel deep architecture that dynamically models high-dimensional urban representations by integrating global spatial dependencies while eliminating the reliance on predefined spatial structures.
  • We introduce two novel modules: a dependency retrieval module that captures urban structures, and a geospatial attention mechanism that prioritizes representations based on spatial proximity and information entropy.
@@
High-Dimensional Representations for Urban Forecasting
Urban forecasting has increasingly leveraged high-dimensional data to address complex challenges. Traditional approaches often rely on low-dimensional numerical data, such as Point of Interest (POI) data(Li et al. 2022), survey data(Roskam and Kunst 2008; DHS et al. 2013), GPS records (Wang et al. 2023; Moreira-Matias et al. 2013; Pappalardo et al. 2013), demographic census, spatial features and so forth, limiting their ability to capture the multifaceted nature of urban dynamics. Recent advances have focused on two primary directions for high-dimensional urban representations.
The first direction employs graph-based methods that embed urban information into representations by constructing predefined graph structures. Graph Attention Networks (GATs) (Zhang, Yu, and Liu 2019) are used to infer POI relationships (Chen et al. 2022b; Li et al. 2020, 2023). Feng and Tassiulas propose an Adaptive Graph Spatial-Temporal Transformer Network to model cross-spatial-temporal correlations. Zheng et al. use a graph multi-attention network to model the impact of the spatio-temporal factors on traffic conditions. Chen et al. propose a spatial-aware attention module based on spatial proximity. Additionally, position embeddings are employed as learnable parameters to provide location information (Xu et al. 2021). These methods rely heavily on predefined spatial structures, such as road networks or adjacency matrices, which limits their applicability in dynamic or data-sparse urban environments.
The second direction focuses on region-based methods that leverage satellite imagery and other high-resolution data to encode urban areas into high-dimensional representations. Tile2Vec (Jean et al. 2018) learns representations from satellite imagery tiles. (Wang et al. 2024) propose a deep hybrid model fusing regional built environment and socio-demographic information into latent representations through multi-task learning. These region-based representation learning methods avoid reliance on predefined graph structures. However, they only leverage local information patches for prediction (Wang et al. 2024) and struggle to incorporate global urban context or dynamically model spatial dependencies. Additionally, region-based representations face variability in effectiveness due to encoding uncertainty, which remains unaddressed. However, information entropy has been shown to evaluate the effectiveness of high-dimensional data (Wu et al. 2024a) and has been used for weighting (Zhu, Tian, and Yan 2020). It offers a method to mitigate these disparities among region-based urban representations, presenting a promising yet underexplored solution.
Urban Spatial Dependency Modeling
@@
Spatial dependency provides key urban structural insights. It has traditionally been addressed through two main approaches for low-dimensional numerical data: feature engineering methods, such as k-nearest neighbors (KNN), and spatial statistical techniques, including spatial error regression, spatial lag models (Anselin 2013), and geographically weighted regression (GWR) (Fotheringham, Charlton, and Brunsdon 1998).
For high-dimensional representations, GATs employ attention mechanisms to aggregate information from neighboring nodes (Feng and Tassiulas 2022; Zheng et al. 2020; Chen et al. 2022b; Xu et al. 2021). However, region-based high-dimensional urban representations often lack explicit graph-like structures, further complicating their ability to capture spatial dependencies. Data-driven methods like sparse linear layers may provide insights to dynamically capture spatial dependency (Li and Moura 2020), yet their applications on region-based spatial representations remain unexplored.
Therefore, to address these challenges, we propose leveraging a sparse matrix-based dependency retrieval mechanism to dynamically model urban structures and integrating geospatial cross-attention weighted by spatial proximity and information entropy to incorporate both local and global information. These techniques aim to overcome the limitations of predefined graph structures and local-only region-based approaches in existing studies.
@@
Gross Domestic Product (GDP) is a critical indicator in urban studies, offering vital insights into the economic health, developmental patterns, and sustainability of cities(Ribeiro et al. 2021; Mahtta et al. 2022; Brilhante and Klaas 2018). Leveraging Nighttime Light (NTL) data to estimate GDP as an economic indicator has become a widely adopted approach (Liu et al. 2021; Sun et al. 2020). There is substantial research indicating a strong association between GDP and urban spatial characteristics (Zhang, He, and Yuan 2020; Zhaoa and Chenb 2023), with numerous studies employing machine learning methods to model and predict regional GDP using structured urban information or urban imagery data(Kopczewska 2022; Chen et al. 2022c). For example, Chen et al. utilize deep learning methods to predict county-level GDP grids. While substantial research has established the relationship between GDP and urban spatial characteristics, the application of high-dimensional urban representations for GDP prediction remains underexplored. Given the critical role of GDP in validating urban prediction models, we selected GDP estimates derived from Nighttime Light (NTL) data as a benchmark for evaluating our approach.
In addition to economic indicators, mobility indicators such as ride-share demand serve as a critical indicator for urban forecasting, and it has significant implications for transportation planning, economic activity, and population behavior (Komanduri et al. 2018; Ma and Hanrahan 2020; Zhang and Zhang 2018). For instance, Shaheen et al. emphasize that ride-share services can influence traffic patterns, reduce private vehicle ownership, and improve accessibility in urban environments. Researchers have developed utilizing deep learning methods to predict ride-share demand from urban spatial information(Mühlematter et al. 2023; Zhao et al. 2022; Tu et al. 2024). Zhao et al. couple graph neural networks (GNNs) with the spatial-temporal influence of the built environment. Zhang and Zhao propose a Clustering-aided Ensemble Model to improve ridesourcing demand forecasting. Similarly, few studies have explored using high-dimensional urban representations to predict ride-share demand. Given the significance of ride-share demand as a key indicator with strong spatial dependencies, it serves as an ideal benchmark for evaluating the efficacy of our proposed method.
@@
GeoTransformer
As illustrated in Figure 1, the whole framework adopts an encoder-decoder structure. The encoder leverages a region-based representation method to process numeric and imagery inputs, representing city regions in a latent space through a mixing operator.
We propose a novel approach GeoTransformer to decode the representations for downstream tasks including GDP and ride-share demand prediction. Our method contains two key innovations: a dependency retrieval module and a geospatial attention module.
@@
As shown in Figure 2, our method adopts a structure akin to the Transformer (Vaswani et al. 2017) decoder. Most relevant regions are retrieved from all urban regions. Then the target and retrieved regions are encoded into representations and then input into the transformer decoder for prediction. The GeoTransformer system $\mathcal{G}$ can be represented as:
$y_{i}=\mathcal{G}(z_{i},\text{Retrieval}(i,Z))$ (1)
in which the input includes the latent representation $z_{i}$ of the target region $i$ and retreived representations $Z_{i}=\text{Retrieval}(i,Z)$ from all regions $Z$, and $y_{i}$ is the predicted output.
Figure 2: The GeoTransformer framework includes a retrieval module and a geospatial attention module that calculates attention scores between the query region and retrieved value regions.
Region-based Urban Representations
To obtain the regional representations, we adopt a mixing operator $\mathcal{M}$ following the work Deep Hybrid Model (Wang et al. 2024), to fuse satellite imagery and sociodemographic information into a high-dimensional latent space, which can be represented as:
$z_{i}=\mathcal{M}(x_{i},I_{i})$ (2)
in which $x_{i}$ and $I_{i}$ represent the sociodemographics and satellite imagery inputs, and $z_{i}$ represents the latent representation of an urban area.
The supervised autoencoder (SAE) is deployed in this framework to learn the representations through optimizing an autoencoder loss and a sociodemographic supervision loss (Figure 3). Each region of a city is then encoded into a latent representation using the encoder of the trained SAE.
@@
Dependency Retrieval Module
To address the challenge of modeling dynamic inter-region interactions for region-based representation methods, we employ a sparse matrix to explicitly represent spatial dependencies. The sparse matrix encodes the influence of one region on another through weights, where only a subset of regions significantly influence the given target region. Building on this, we propose a dependency retrieval mechanism that dynamically identifies the most relevant regions around the target region as the GeoTransformer model inputs. Specifically, for each target region, the mechanism first identifies the $k_{1}$-nearest neighbors based on spatial proximity. Then, from these $k_{1}$ candidates, it selects the top-$k_{2}$ regions with the highest weights in the sparse matrix. The retrieval mechanism is defined as:
$Z_{i}=\{z_{j}\mid j\in\mathrm{arg\,top}_{k_{2}}(S[i,\mathcal{N}_{k_{1}}(i)])\},$ (3)
where $Z_{i}$ denotes the set of high-dimensional representations $z_{j}\in\mathbb{R}^{d}$ for the $k_{2}$ most relevant regions selected from the $k_{1}$-nearest neighbors of the target region $i$. Here, $\mathcal{N}_{k_{1}}(i)$ represents the indices of the $k_{1}$-nearest neighbors of region $i$, identified based on spatial proximity. $S[i,j]$ is the weight from the sparse matrix $S$, which quantifies the dependency of region $i$ on region $j$.
The sparse matrix $S$ is constructed using Lasso regression, which enforces sparsity to identify the most important inter-region dependencies. For a given target region $i$, its high-dimensional representation $z_{i}\in\mathbb{R}^{d}$ is treated as the target variable, while the representations of all other regions $X$ serve as input features. The optimization objective for Lasso is defined as:
$\min_{\beta}\|z_{i}-X\beta\|_{2}^{2}+\alpha\|\beta\|_{1},$ (4)
where $X\in\mathbb{R}^{(n-1)\times d}$ denotes the matrix of feature representations for all other regions except $i$. The sparse weight vector $\beta\in\mathbb{R}^{n-1}$ captures the influence of other regions on region $i$, while $\alpha$ is the regularization strength that controls the degree of sparsity in the resulting weights. The Lasso regularization term $\|\beta\|_{1}$ forces most weights to zero, retaining only the strongest dependencies. The resulting sparse weight vector $\beta$ is then used to populate the corresponding row $S[i,:]$ in the sparse matrix $S$.
This dependency retrieval module enables region-based representations to effectively model spatial dependency without predefined urban structure information, providing a foundation for attention computing in the GeoTransformer model.
Geospatial Attention Module
We introduce a geospatial attention mechanism to leverage the global urban information for prediction. As shown in Figure 2, the model first calculates cross-attention between the target and retrieved representations, with the attention scores weighted by spatial proximity and information entropy. The attention calculation is represented as:
$\text{GeoAtt}(Q,K,V)=(\lambda W_{s}+(1-\lambda)W_{e})\odot\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (5)
@@
where the query matrix $Q$ contains the representation of the query region $z_{i}$ for which the target variable is being predicted, while the value matrix $V$ consists of $z_{i}$ itself and the retrieved representations as value regions. $d_{k}$ denotes the dimension of the key vectors. Intuitively, as the similarity between two city regions is not directly related to prediction variables, we set the key matrix $K$ as trainable weights rather than being identical to $V$. $W_{s}$ and $W_{e}$ are the spatial proximity and information entropy weighting factors respectively. Their relative importance is controlled by a hyperparameter $\lambda\in[0,1]$.
Spatial Proximity Weighting
@@
Our model provides spatial proximity weighting as a way to provide location information. Based on the assumption of Tobler’s First Law of Geography (Tobler 1970), we assigned regions closer to the query region with higher weights. We leverage a linear weighting method, scaling the distances within 0 to 1, inversely transforming shorter distances into higher weights:
$W_{S_{j}}=1-\frac{d_{j}}{\text{max}(d)}$ (6)
@@
Information Entropy Weighting
As a way to evaluate the effectiveness of high-dimensional data, we assign higher weights for representations with higher information entropy. Regions with higher entropy are believed to contain richer and more complex urban information, thereby playing a more significant role in prediction. The calculation can be presented as:
$W_{E_{j}}=\frac{H_{j}}{\max(H_{1},H_{2},\dots,H_{n})}$ (7)
where $H_{i}$ is the Shannon entropy of the latent representation $z_{i}$, and $\max(H_{1},H_{2},\dots,H_{n})$ is the maximum entropy across retrieved tiles. Since the the latent representation is a high-dimensional vector $z_{i}$, the entropy $H_{i}$ for each individual latent representation cab be represented as:
${H}_{j}=-\sum_{i=1}^{d}p_{ji}\log(p_{ji})$ (8)
where $\textbf{where }\quad\text{p}_{ij}=\frac{e^{z_{ij}}}{\sum_{k=1}^{d}e^{z_{ik}}}$ (9)
Here, $z_{ji}$ is the value of the $j-{th}$ feature for latent representation $z_{i}$, and $p_{ij}$ is the normalized probability.
Adhering to the established multi-head attention mechanism (Vaswani et al. 2017), our model employs multiple attention heads to capture various spatial relationships concurrently. The calculation can be represented as:
@@
where $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$ and $W^{O}$ are the projection matrices. The amount of heads and layers of the decoder module is also adaptive. At last, the outputs of the decoder layers are passed through a fully connected linear layer.
The GeoTransformer framework is designed to overcome limitations in region-based urban representation methods, capturing spatial dependency dynamically and leveraging global urban information for prediction. It is also adaptable to various urban prediction tasks due to the flexibility of the output layers. Overall, our GeoTransformer represents a novel approach to geospatial analysis and urban planning applications.
Panel 1 Deep Hybrid Model Framework (Latent Representations + Regression)
Model 1: DHM (linear) representation input (4096 dim), Linear Regression
Model 2: DHM (polynomial) representation input (reduced dim from 4096 to 128 using PCA), Polynomial Regression (poly degree = 2, $l_{2}$ regularization $\lambda$ = 0.5)
Panel 2 Latent Representations + Graph Attention Networksl
Model 3: GAT (grid) representation input (4096 dim), layers = 4
Model 4: GAT (sparse) representation input (4096 dim), layers = 4, retrieval size = 81
Panel 3 Latent Representations + GeoTransformer (Ours)
Model 5: GeoTransformer representation input (4096 dim), heads = 16, layers = 4, KNN=121, retrieval size = 81
@@
Panel 1 Deep Hybrid Model Framework (Latent Representations + Regression)
1 DHM (linear) 3.56/3.23 1.84/1.75 0.151/0.142 9.03/7.28 9.36/8.90 0.568/0.601
2 DHM (polynomial) 2.19/3.25 1.17/1.46 0.429/0.118 12.8/19.5 5.05/6.00 0.494/0.335
Panel 2 Latent Representations + Graph Attention Networks
3 GAT (grid) 0.469/0.726 0.476/0.602 0.866/0.805 7.21/9.51 2.01/2.35 0.750/0.575
4 GAT (sparse) 0.592/0.691 0.547/0.561 0.852/0.812 6.22/5.47 1.66/1.83 0.781/0.794
Panel 3 Latent Representations + GeoTransformer (Ours)
5 GeoTransformer 0.079/0.216 0.207/0.343 0.964/0.933 0.217/2.94 0.684/1.02 0.920/0.901
@@
In this section, we provide our method’s experiments in GDP and ride-share demand prediction and compare them with baseline methods. Ablation experiments are conducted to demonstrate the effectiveness of the dependency retrieval module and two weighting methods in the geospatial attention module.